OpenStudy (anonymous):

arc length of y=(1/4)x^2-(1/2)ln(x) from x=1 to 2

OpenStudy (anonymous):

these problems are always cooked up to be doable did you find \(y'\)?

OpenStudy (anonymous):

i think it is \[y'=\frac{1}{2}x-\frac{1}{2x}\]

OpenStudy (anonymous):

then you have to square it

OpenStudy (anonymous):

something like \[\frac{(x^2-1)^2}{4 x^2}\]

OpenStudy (anonymous):

then you have to add one to it

OpenStudy (anonymous):

and that is where the miracle occurs you get \[\frac{(x^2+1)^2}{4 x^2}\] which is a perfect square

OpenStudy (anonymous):

good thing since now you have to take the square root of it that gives you \[\int_1^2\frac{x^2+1}{2x}dx\]

OpenStudy (anonymous):

imagine how much work went in to making this problem work out so nicely i tried it once took me over half an hour to make one

OpenStudy (kanwal32):

@satellite73 i did the same way as u said to calculate Arithmetic mean(before asking)